# Self-Quenching Effect of the Decay of Localized Surface Plasmons: Classical and Quantum Perspectives

### v1

@@ 5 The Dual Functionality of Plasmonic Quasi-Particles
In contrast, the excited plasmonic nanoparticle plays a dual role, acting as a source of plasmonic cavity and the TM emitter embedded in a cavity of confined longitudinal near-field modes. The same physical structure allows the collective surface charge oscillations for a certain time in the femtosecond range (Fig. 1a)) after the LSP excitation. These oscillations are a source of the locally confined longitudinal fields, which leak from the surface, and decay as $1/r^{2}$ (see Section 2). The longitudinal evanescent field modes coupled to the oscillating surface charges do not radiate energy away. Instead, they act as a storage of reactive energy, while the tangential field components decay as $1/r$, carrying part of the energy to the far field region.
@@
  • The emitter and the cavity share identical spectral parameters, such that $\omega_{l}^{c}=\omega_{l}$ and $\Gamma_{l}^{c}=\Gamma_{l}=\Gamma_{l}^{reff}+\Gamma^{nr}$.
@@ 5.1 Quality Factor of Plasmonic Cavity
The quality factor of the plasmonic cavity $Q_{l}(R)$, [derom2012resonance, kristensen2014modes, stockman2018roadmap, kolwas2020impact] defined as the number of oscillations required for the energy of a freely oscillating system to fall off to $e^{-2\pi}$ of its original energy, is a function of the radius $R$ of MNP:
@@
where $\Gamma_{l}^{c}$ is composed of the radiative $\Gamma_{l}^{reff}$ and nonratiative $\Gamma^{nr}$ contributions (Eq (7)).
The quality factor thus encapsulates both the intrinsic resonant confinement of the EM field (via $\omega_{l}$) and the loss mechanisms that limit the ability of the cavity to store the total energy of LSP excitation. It plays a central role in determining LDOS, which in turn controls the emission dynamics of the plasmonic emitter.
5.2 Density of States of Plasmonic Cavity Created by Plasmonic Quasiparticle
@@
where $\Delta\omega_{l}^{c}$ is the full width at half maximum (FWHM) of the cavity resonance, linked to the total loss rate $\Delta\omega_{l}^{c}=\Gamma_{l}^{c}=\Gamma_{l}(R)$ (Eqs. (7), (22) of the mode and $Q_{l}$ is the quality factor.
@@
Conversely, traditional CQED and PCQED systems use an external, independent resonators. Within the PQP framework, the correspondence between the emitter and cavity parameters necessitates a novel, self-consistent formulation of the spontaneous emission decay from the confined plasmonic systems.
5.3 Extension of Fermi’s Golden Rule to PQP Systems
@@
In contrast, a plasmonic quasi-particle (PQP) forms a self-consistent system where both the emitter and the cavity originate from the same physical structure and share the same spectral parameters. The emission occurs into a structured, near-field environment that is dynamically coupled to the emitter.
@@
where: $\Gamma_{l}^{p}=2\Gamma_{l}^{ref}$ is the spontaneous emission rate, and $\Gamma_{l}^{c}=\Gamma_{l}^{reff}+\Gamma^{nr}$ is the total linewidth of the PQP cavity. The matrix element $|\langle l|\hat{V}|0\rangle|^{2}$ is renormalized by a mode confinement factor $\beta_{l}=V_{l}/V_{l}^{eff}$.
In the collisionless limit ($\Gamma^{nr}=0$), Eq. (25) simplifies to:
@@
$\Gamma_{l}^{p}=2\Gamma_{l}^{reff}=\frac{2\Gamma_{l}^{r}\Gamma_{l}^{c\,r}}{\Gamma_{l}^{c}}=\frac{2(\Gamma_{l}^{r})^{2}}{\Gamma_{l}^{reff}+\Gamma^{nr}},$ (28)
which yields the self-consistent radiative emission rate:
$\Gamma_{l}^{reff}=\frac{1}{2}\left(\sqrt{4(\Gamma_{l}^{r})^{2}+(\Gamma^{nr})^{2}}-\Gamma^{nr}\right)\leq\Gamma_{l}^{r}.$ (29)
This expression reveals a self-quenching effect, in which a nonradiative decay channel attenuates radiative emission as a consequence of the dynamic coupling between the emitter and the cavity. This mechanism is fundamentally different from classical Purcell suppression, since it arises from the dissipative characteristics of the emitter’s immediate self-created EM environment.
Nonetheless, it can be viewed as analogous to fluorescence quenching in conventional PQED architectures (Section 4.2), which occurs when an atomic or molecular emitter is placed in close proximity to a plasmonic nanoparticle [lakowicz2005radiative, anger2006enhancement, bharadwaj2006nanoplasmonic, guzatov2012plasmonic, olejnik2012plasmonic, rasskazov2018plasmon, kongsuwan2018suppressed, hugall2018plasmonic, hildebrandt2023plasmonic].
@@
If $\Gamma^{nr}\ll\Gamma_{l}^{r}$, then $\Gamma_{l}^{reff}\approx\Gamma_{l}^{r}$;
If $\Gamma^{nr}\gg\Gamma_{l}^{r}$, then $\Gamma_{l}^{reff}\ll\Gamma_{l}^{r}$.
@@
$\Gamma_{l}(R)=\Gamma_{l}^{reff}(R)+\Gamma^{nr}(R)\leq\Gamma_{l}^{r}(R)+\Gamma^{nr}(R).$ (30)
@@
The structure of Eq. 29 explicitly illustrates the non-additive character of the decay channels. The intrinsic nonradiative losses in metal NPs do not merely contribute linearly to the LSP damping. Instead, they can dynamically suppress the total decay rate within certain regions of the parameter space.
The reduction of the total damping rate $\Gamma_{l}(R)$ is consistent with our predictions based on classical electrodynamics (Eq. (7)). The quenching effect manifests in low-radiative modes within the size and multipolarity ranges for which $0<\Gamma_{l}^{r}(R)\leq\Gamma_{l}^{nr}(R)$ (Fig. 2b).
While the self-quenching effect described by Eq. 29 is consistent with the qualitative trends observed in our classical results (Section 2), a quantitative validation requires separating the radiative component $\Gamma_{l}^{r}(R)$ from the total damping rate $\Gamma_{l}(R)$. This decomposition was not part of our previous classical analysis, which focused on total damping. Additionally, explicit parametrization of interband transitions, which dominate $\Gamma^{nr}(R)$ at small radii, is necessary but currently incomplete. A detailed classical study addressing both aspects is underway and will be reported separately.
5.4 Correspondence to the experimental observations
@@
Experimental investigations of localized surface plasmon (LSP) damping in noble-metal nanoparticles conventionally represent the total damping rate as the linear superposition of independent radiative and nonradiative contributions. Nonetheless, several studies (e.g. [sonnichsen2002drastic, moustafa2023bandwidth, kolwas2013damping, ginzburg2012non] have reported anomalous behavior that departs from the standard assumptions in specific systems and experimental conditions.
@@
In [sonnichsen2002drastic], the authors study plasmon damping rates in single gold nanoparticles using dark-field spectroscopy to record light-scattering spectra as a function of particle size and shape. The study demonstrates that in the spherical NPs, the homogeneous broadening (dephasing rate) of the plasmon resonance increases with particle size due to enhanced radiation damping, as expected. In contrast, in nanorods, the dephasing rate of a low radiative dipole LSP resonance drastically decreases with increasing length. The authors attribute this unexpected effect to suppressed interband damping, which contributes to nonradiative damping. Their supporting quasistatic calculations neglect radiation damping.
@@
The results for the dipole mode of nanospheres remain in accordance with our expectations in the studied range of sizes $2R=$ 20,…, 150 nm, as illustrated in Fig. 2b)) (black line).
@@
Nanorods, in contrast, offer more favorable conditions for observing the self-quenching effect. The long-axis mode behaves nearly as an ideal dipolar emitter with low (suppressed) radiative losses, even for larger sizes. Therefore, the condition $\Gamma^{nr}>\Gamma_{l}^{r}$ is fulfilled. According to Eq. 29), in such a range of parameters, one can expect suppression of the radiative damping with the increasing NP size, as illustrated for higher-order modes in spherical NPs (Fig. 2b)).
@@
More direct evidence of the coupling between radiative and nonradiative channels was provided by recent TR-PEEM (Time-Resolved PhotoEmission Electron Microscopy) experiments on gold nanorods [qin2023coaction]. The authors measured lifetimes of bright (superradiant) and dark (subradiant) modes as a function of nanorod length. Surprisingly, bright modes were found to have lifetimes equal to or longer than those of subradiant modes, contradicting the expectation that radiative losses shorten the lifetimes of bright modes: $\Gamma^{super}(\Gamma^{r},\Gamma^{nr})\leq\Gamma^{subr}(\Gamma^{nr})$, despite the presence of radiative damping in the superradiant case. This confirms the co-dependence of decay channels Eq. 29).
This behavior directly supports our prediction (Eq. 29). It demonstrates the non-additivity of decay channels. The presence of inherent nonradiative losses in metal NPs does not simply add to the LSP damping. Instead, it can dynamically suppress the total decay rates in some parameter ranges.
@@ 6 Conclusions
This work presents a unified, quantum-informed framework in which decaying localized surface plasmon (LSP) excitations are conceptualized as plasmonic quasi-particles - quantized, coherent many-electron emitters that radiate energy into their own structured near-field TM-polarized electromagnetic environment. The excited plasmonic emitter and its self-generated EM environment form the cavity of TM confined field modes, sharing the same transient dynamics as the source fields of LSP. Therefore, both the decay rates of surface TM-polarized fields and cavity fields are composed of radiative and nonradiative contributions.
@@
An intrinsic suppression of the total damping rate of radiative modes due to the self-quenching effect is predicted. This effect does not arise from an externally altered EM environment as in the traditional Purcell effect. Rather, it occurs due to the loss of part of the excitation energy via collisional, ohmic losses in a metal NP, resulting in heat generation. Consequently, the emission characteristics of the plasmonic system are modified self-consistently, as the effective LDOS, dependent on the quality factor of the self-created plasmonic cavity, is influenced by the nonradiative losses process, which introduces a co-dependence of radiative and non-radiative decay channels.
Thus, the conventional view of LSP decay as a sum of independent radiative and nonradiative channels is no longer sufficient in general. Instead, radiative emission is inherently reduced due to the self-quenching effect.
@@
These result aligns with our description of the transient LSP decay in classical frames of Maxwell electrodynamics and help explain experimental observations of anomalously long lifetimes in nanorods [sonnichsen2002drastic, qin2023coaction]. Classically derived cavity parameters, obtained from the dispersion relation of quasi-normal transverse-magnetic (TM) multipolar modes (retardation effects included), form the foundation of our model. Their reinterpretation within the self-consistent quantum framework offers a physically transparent bridge between Maxwell and quantum descriptions of LSP dynamics.
@@
Beyond its implications for emission control, the model highlights the bosonic nature of PQPs: coherent excitations of many-electron systems that behave collectively as quantized states. These systems can be viewed as micro- and macroscopic quantum structures, sharing characteristics related to the phenomenon of macroscopic quantum coherence. These localized surface-bound bosonic modes form without nonlinear thresholds and can exist at room temperature. Unlike traditional Bose–Einstein condensates (BECs), which require ultralow temperatures and massive particles, PQPs provide a platform for studying room-temperature bosonic phenomena in three-dimensional systems. This insight aligns with recent demonstrations of polariton and quasiparticle condensation under ambient conditions [lagoudakis2010coherent, kkedziora2024predesigned], further validating the relevance of the PQP model.
@@
The self-consistent, quantum-local framework for a deeper understanding of the decay process in coherent many-electron micro- and macro-systems presented here can possibly serve in the development of decay engineering, controlled light–matter interactions, and the development of future quantum technologies based on highly dissipative yet coherent plasmonic systems.
7 Funding Declarations
@@
8 Availability of data and materials
All data supporting the findings of this study are either presented within the manuscript or based on previously published results, as referenced in [kolwas2013damping]: K. Kolwas and A. Derkachova, "Damping rates of surface plasmons for particles of size from nano- to micrometers; reduction of the nonradiative decay," JQSRT, 114 (2013), pp. 45–55. No new datasets were generated or analyzed specifically for this work.

### v2

@@ 4 Dual Functionality of Plasmonic Quasi-Particles
In contrast, the excited plasmonic nanoparticle plays a dual role, acting as a source of plasmonic cavity and the TM emitter embedded in a cavity of confined longitudinal near-field modes. The same physical structure allows the collective surface charge oscillations for a certain time in the femtosecond range (Fig. 1a)) after the LSP excitation. These oscillations are a source of the locally confined longitudinal fields, which leak from the surface and decay as $1/r^{2}$. The longitudinal evanescent field modes coupled to the oscillating surface charges do not radiate energy away. Instead, they act as a storage of reactive energy, while the tangential field components decay as $1/r$, carrying part of the energy to the far field region.
@@
  • The emitter and the cavity share identical spectral parameters, such that $\omega_{l}^{c}=\omega_{l}$ and $\Gamma_{l}^{c}=\Gamma_{l}=\Gamma_{l}^{reff}+\Gamma_{l}^{nr}$.
@@ 4.1 Quality Factor of Plasmonic Cavity
The quality factor of the plasmonic cavity $Q_{l}(R)$, 11, 34, 59, 30 defined as the number of oscillations required for the energy of a freely oscillating system to fall off to $e^{-2\pi}$ of its original energy, is a function of the radius $R$ of MNP:
@@
where $\Gamma_{l}^{c}$ is composed of the radiative and nonratiative contributions (Eq (7)).
The quality factor thus encapsulates both the intrinsic resonant confinement of the EM field (via $\omega_{l}^{c}$) and the loss mechanisms that limit the ability of the cavity to store the total energy of LSP excitation. It plays a central role in determining LDOS, which in turn controls the emission dynamics of the plasmonic emitter.
4.2 Density of States Available for Plasmonic Emitter
@@
where $\Delta\omega_{l}^{c}$ is the full width at half maximum (FWHM) of the cavity resonance, linked to the total loss rate $\Delta\omega_{l}^{c}=\Gamma_{l}^{c}=\Gamma_{l}$ (Eqs. (7), (22) of the mode and $Q_{l}$ is the quality factor.
@@
Conversely, traditional CQED and PCQED systems use external, independent resonators. Within the PQP framework, the correspondence between the emitter and cavity parameters necessitates a novel, self-consistent formulation of the spontaneous emission decay from the confined plasmonic systems.
4.3 Extension of Fermi’s Golden Rule to POP Systems
@@
By contrast, a plasmonic quasi-particle (PQP) constitutes a self-consistent, strongly coupled system in which both the emitter and the EM environment arise from the same underlying physical structure and are characterized by identical spectral parameters. In this case, radiation is emitted into a spatially structured, near-field EM environment that remains dynamically coupled to the emitter.
@@
where $\Gamma_{l}^{p}=2\Gamma_{l}^{r\,eff}$ denotes the spontaneous emission rate of the plasmonic quasi-particle, and $\Gamma_{l}^{c}=\Gamma_{l}^{r\,eff}+\Delta_{l}+\Gamma^{nr}$ is the total linewidth of the LSP resonance (Eq. 7) and therefore that of the PQP cavity (Eq. 22), with parameters adjusted to the damping rates of the emitter. The reduction term $\Delta_{l}$, which arises from the nonorthogonality of the eigenstates of the open system has been incorporated in $\Gamma_{l}^{nr}$. The matrix element $|\langle l|\hat{V}|0\rangle|^{2}$ is renormalized by a mode confinement factor $\beta_{l}=V_{l}/V_{l}^{eff}$.
The collisionless limit ($\Gamma_{l}^{nr}=0$), Eq. (25), allows us to connect the radiative rate $\Gamma_{l}^{r}$ to the perturbation strength between the final and initial states:
@@
$\Gamma_{l}^{p}=2\Gamma_{l}^{reff}=\frac{2\Gamma_{l}^{r}\Gamma_{l}^{c\,r}}{\Gamma_{l}^{c}}=\frac{2(\Gamma_{l}^{r})^{2}}{\Gamma_{l}^{reff}+\Gamma_{l}^{nr}}.$ (28)
Solving Eq. (28) self-consistently yields the effective radiative decay rate of the plasmonic quasi-particle. The resulting analytical expression constitutes the central result of the present model, as it explicitly reveals the feedback between intrinsic nonradiative dissipation and the radiative emission process:
$\Gamma_{l}^{reff}=\frac{1}{2}\left(\sqrt{4(\Gamma_{l}^{r})^{2}+(\Gamma_{l}^{nr})^{2}}-\Gamma_{l}^{nr}\right)\leq\Gamma_{l}^{r}.$ (29)
In the limiting case $\Gamma_{l}^{nr}\rightarrow 0$, Eq. (29) correctly reduces to $\Gamma_{l}^{r\,eff}=\Gamma_{l}^{r}$, recovering the conventional radiative decay rate.
Expression 29 reveals a self-quenching effect, in which a nonradiative decay channel attenuates radiative emission as a consequence of the dynamic coupling between the emitter and the cavity. This mechanism is fundamentally different from classical Purcell suppression, since it arises from the dissipative characteristics of the emitter’s immediate self-created EM environment.
Nonetheless, it can be viewed as analogous to fluorescence quenching in conventional PQED architectures, which occurs when an atomic or molecular emitter is placed in close proximity to a plasmonic nanoparticle 37, 2, 5, 18, 47, 51, 33, 21, 20.
@@
If $\Gamma_{l}^{nr}\ll\Gamma_{l}^{r}$, then $\Gamma_{l}^{reff}\approx\Gamma_{l}^{r}$;
If $\Gamma_{l}^{nr}\gg\Gamma_{l}^{r}$, then $\Gamma_{l}^{reff}\ll\Gamma_{l}^{r}$.
@@
$\Gamma_{l}=\Gamma_{l}^{r\,eff}+\Gamma_{l}^{nr}\leq\Gamma_{l}^{r}(R)+\Gamma_{l}^{nr}.$ (30)
@@
The structure of Eq. (29) explicitly illustrates the non-additive character of the decay channels. The nonradiative, collisional losses in metal NPs do not merely contribute linearly to the LSP damping. Instead, they can dynamically suppress the total decay rate within certain regions of the parameter space.
Although the self-quenching effect described by Eq. (29) supports the central claim of this paper that plasmon damping cannot in general be treated as a simple sum of independent decay channels (Eq. (7) and Fig. 2), a quantitative evaluation of the effect versus size requires separating the radius dependence of the radiative contribution $\Gamma_{l}^{r}(R)$ from the total damping rate $\Gamma_{l}(R)$. Such a decomposition was not performed in Ref 28. In addition, a complete parametrization of interband transitions, which contribute to nonradiative damping rates $\gamma_{l}^{nr}(R)$ in some size ranges, remains incomplete. A systematic study addressing these aspects is currently in progress and will be reported separately.
4.4 Comments on Some Experimental Results in Nanorods
@@
Experimental investigations of localized surface plasmon (LSP) damping in noble-metal nanoparticles conventionally represent the total damping rate as the linear superposition of independent radiative and nonradiative contributions. Nonetheless, several studies (e.g. 58, 44, 28, 16 have reported anomalous behavior that departs from the standard assumptions in specific systems and experimental conditions.
@@
In 58, the authors study plasmon damping rates in single gold nanoparticles using dark-field spectroscopy to record light-scattering spectra as a function of particle size and shape. The study demonstrates that in the spherical NPs, the homogeneous broadening (dephasing rate) of the dipole plasmon resonance increases with particle size due to enhanced radiation damping, as expected. In contrast, in nanorods, the total dephasing rate of a low-radiative LSP dipole resonance decreases drastically with increasing length. The authors attribute this unexpected effect to suppressed interband damping, which contributes to nonradiative damping. Their supporting quasistatic calculations neglect radiation damping.
@@
The results for the dipole mode of nanospheres remain in accordance with our expectations in the studied range of sizes $2R=$ 20,…, 150 nm, as illustrated in Fig. 2b) (black line).
@@
Nanorods, in contrast, offer more favorable conditions for observing the coupling between the radiative and nonradiative channels, thereby reducing the total damping rate. The long-axis mode behaves nearly as an ideal dipolar emitter with low radiative losses, even for larger sizes. Therefore, the condition $\Gamma_{1}^{nr}\gtrsim\Gamma_{1}^{r}$ is fulfilled over a wider range of sizes than in the case of spherical MNPs. According to Eqs. (29) and Eqs. (30) in such a range of parameters, one can expect suppression of the radiative (and the total damping) rate with the increasing NP size.
@@
More direct evidence of the coupling between radiative and nonradiative channels was provided by recent TR-PEEM (Time-Resolved PhotoEmission Electron Microscopy) experiments on gold nanorods 50. The authors measured lifetimes of bright (superradiant) and dark (subradiant) modes as a function of nanorod length. Surprisingly, bright modes were found to have lifetimes equal to or longer than those of subradiant modes, contradicting the expectation that radiative losses shorten the lifetimes of bright modes: $\Gamma^{super}(\Gamma^{r},\Gamma^{nr})\leq\Gamma^{subr}(\Gamma^{nr})$, despite the presence of radiative damping in the superradiant case. This confirms the co-dependence of decay channels Eq. 29).
Such behavior provides qualitative support for our prediction, evidencing the non-additive character of the decay channels that results in a reduction of the experimentally observed total damping rates in metallic nanoparticles. In plasmonic systems, collisional, heat-generating losses can dynamically suppress the overall decay rates within specific regions of parameter space.
@@ 5 Conclusions
This work introduces a unified, quantum-informed framework in which decaying localized surface plasmon (LSP) excitations are treated as plasmonic quasiparticles: quantized, coherent many-electron emitters that radiate energy into their own structured electromagnetic near-field environment of longitudinal fields. In this description, the plasmonic excitation simultaneously acts as both the emitter and the source of the confined electromagnetic modes forming the effective cavity. Unlike conventional cavity quantum electrodynamics (CQED), where the emitter and cavity are treated as independent subsystems, the present framework considers them as dynamically coupled parts of the same physical system. The implications of this intrinsic coupling for the damping dynamics of coherent electron oscillations constitute the central conceptual aspect of the present study.
@@
An intrinsic suppression of the total damping rate of radiative modes due to the self-quenching effect is predicted. Unlike in the traditional Purcell effect, this modification of the damping rate does not arise from an externally altered EM environment. Rather, it results from the loss of a portion of the excitation energy via collisional, ohmic processes in a metal nanoparticle, leading to heat generation. Consequently, the emission characteristics of the plasmonic system are modified self-consistently: the effective LDOS, which depends on the quality factor of the self-created plasmonic cavity, is itself influenced by nonradiative losses, introducing a strict co-dependence of the radiative and nonradiative decay channels.
The original contribution of this work lies in the formulation of a self-consistent extension of Fermi’s Golden Rule for plasmonic quasi-particles, in which the radiative decay rate depends explicitly on the total cavity linewidth entering the LDOS. This feedback mechanism leads to an analytical expression for the effective radiative rate (Eq. 29), revealing a non-additive coupling between decay channels and predicting a size- and mode-dependent self-quenching effect. Unlike conventional phenomenological approaches that decompose the total linewidth into independent radiative and nonradiative terms, the present framework demonstrates that intrinsic dissipation modifies the emission process itself through the shared spectral parameters of the emitter and its self-generated cavity. Thus, the conventional description of LSP decay as a simple sum of independent radiative and nonradiative channels fails to fully capture the dynamics of plasmonic systems in which the emitter and cavity are intrinsically coupled.
These results are consistent with the classical description of transient LSP decay within Maxwell electrodynamics and help explain experimental observations of anomalously long lifetimes in nanorods 58, 50.
@@
Classically derived cavity parameters, obtained from the dispersion relation of quasi-normal transverse-magnetic (TM) multipolar modes—including surface scattering and retardation effects—form the physical basis of our model. This framework remains valid over a broad size range, from individual nanometer-scale structures up to several hundreds of nanometers, and extends well beyond the dipole approximation. The reinterpretation of these classical parameters within a self-consistent quantum framework provides a physically transparent bridge between the Maxwell and quantum descriptions of LSP dynamics.
@@
Beyond its implications for emission control, the model highlights the bosonic nature of PQPs: coherent excitations of many-electron systems that behave collectively as a single quantized state. These systems can be viewed as micro- and macroscopic quantum structures sharing characteristics related to macroscopic quantum coherence. These localized surface-bound bosonic modes form without nonlinear thresholds and can exist at room temperature. Unlike traditional Bose–Einstein condensates (BECs), which require ultralow temperatures and massive particles, PQPs provide a platform for studying room-temperature bosonic phenomena in three-dimensional systems. This insight aligns with recent demonstrations of polariton and quasiparticle condensation under ambient conditions 36, 25, further supporting the broader relevance of the PQP framework.
@@
The self-consistent, quantum-local approach presented here contributes to a deeper understanding of decay processes in coherent many-electron micro- and macro-systems. Ultimately, it may serve in the development of decay engineering strategies, controlled light–matter interactions, and future quantum technologies based on highly dissipative yet coherent plasmonic systems.
{acknowledgement}